# Riemannian Consistency Model

### v1

@@ Abstract
Consistency models are a class of generative models that enable few-step generation for diffusion and flow matching models. While consistency models have achieved promising results on Euclidean domains like images, their applications to Riemannian manifolds remain challenging due to the curved geometry. In this work, we propose the Riemannian Consistency Model (RCM), which, for the first time, enables few-step consistency modeling while respecting the intrinsic manifold constraint imposed by the Riemannian geometry. Leveraging the covariant derivative and exponential-map-based parameterization, we derive the closed-form solutions for both discrete- and continuous-time training objectives for RCM. We then demonstrate theoretical equivalence between the two variants of RCM: Riemannian consistency distillation (RCD) that relies on a teacher model to approximate the marginal vector field, and Riemannian consistency training (RCT) that utilizes the conditional vector field for training. We further propose a simplified training objective that eliminates the need for the complicated differential calculation. Finally, we provide a unique kinematics perspective for interpreting the RCM objective, offering new theoretical angles. Through extensive experiments, we manifest the superior generative quality of RCM in few-step generation on various non-Euclidean manifolds, including flat-tori, spheres, and the 3D rotation group SO(3).
@@ 1 Introduction
Besides Euclidean domains like images, generative models on various Riemannian manifolds have a potentially profound impact in scientific domains, including protein generation [55, 3], peptide design [26, 25], robotics [5], and geoscience [41]. For example, the generative modeling of an amino acid requires descriptors of its position, orientation, and torsion angles. While the position is Euclidean, the orientation lies in the 3D rotation group $\mathrm{SO}(3)$ and the torsion angle lies in the flat torus where a translation by $2\pi$ will result in the same angle. While the current Euclidean consistency model has achieved superior performance on images where the distance between two predictions is measured with the standard Euclidean norm, its counterpart in Riemannian manifolds, the Riemannian consistency model (RCM), poses additional challenges and remains largely unexplored. Specifically, the intrinsically curved manifold requires the consistency parameterizations to lie on the manifold. In this way, simple linear interpolation like Lu and Song [33], Yang et al. [54] is not always feasible. Furthermore, such a manifold constraint will impose an additional constraint on the vector field when the consistency loss is enforced, that is, the vector field at different points should also lie in their corresponding tangent spaces, necessitating additional corrections in the time derivative.
@@ B.2 Spherical Manifold
$\exp_{x}(u+\mathrm{d}u)$ $\approx J_{R}^{-1}(y)J_{R}(u)\mathrm{d}u,y=\exp_{x}u,$ (55)
$\exp_{x+\mathrm{d}x}(u)$ $\approx J_{L}^{-1}(y)J_{L}(u)\mathrm{d}x,y=\exp_{x}u.$ (56)
Combining the above approximations, we have
@@
Similarly, we can calculate $\mathrm{d}(\exp u)_{x}$ using the above approximations with the additional covariant derivative term as:
$\mathrm{d}(\exp u)_{x}(v)=R_{u}^{-1}J_{R}(x)v-\frac{1}{2}J_{R}(u)[v,u],$ (58)
where $R_{u}$ denotes the rotation by the rotation vector $u$, and the second term again arises from the covariant derivative. The time derivative of the model using the Jacobian-vector product (JVP) should also be modified as:
@@
where the additional left inverse Jacobian comes from the expansion of the reference vector field $\dot{x}$ using the right expansion BCH formula in Eq.52.
@@ C.3 Training Technique
To further stabilize the gradient variance in consistency models, we build upon the concept of tangent normalization proposed in Lu and Song [33]. We observed that when the tangent norm is small, standard normalization preserves only directional information, which can impede further optimization. Therefore, drawing inspiration from gradient clipping, we introduce a ’tangent clipping’ mechanism, detailed below.
@@
where $M$ represents the maximum tangent norm, which we set to 1 in our experiments. This approach is designed to maintain low gradient variance without inappropriately rescaling the tangent when its norm is small.

### v2

@@ Abstract
Consistency models are a class of generative models that enable few-step generation for diffusion and flow matching models. While consistency models have achieved promising results on Euclidean domains like images, their applications to Riemannian manifolds remain challenging due to the curved geometry. In this work, we propose the Riemannian Consistency Model (RCM), which, for the first time, enables few-step consistency modeling while respecting the intrinsic manifold constraint imposed by the Riemannian geometry. Leveraging the covariant derivative and exponential-map-based parameterization, we derive the closed-form solutions for both discrete- and continuous-time training objectives for RCM. We then demonstrate theoretical equivalence between the two variants of RCM: Riemannian consistency distillation (RCD) that relies on a teacher model to approximate the marginal vector field, and Riemannian consistency training (RCT) that utilizes the conditional vector field for training. We further propose a simplified training objective that eliminates the need for the complicated differential calculation. Finally, we provide a unique kinematics perspective for interpreting the RCM objective, offering new theoretical angles. Through extensive experiments, we manifest the superior generative quality of RCM in few-step generation on various non-Euclidean manifolds, including flat-tori, spheres, and the 3D rotation group SO(3), spanning a variety of crucial real-world applications such as RNA and protein generation. Our code is available at https://github.com/ccr-cheng/riemannian-consistency-model.
@@ 1 Introduction
Besides Euclidean domains like images, generative models on various Riemannian manifolds have a potentially profound impact in scientific domains, including protein generation [55, 3], peptide design [26, 25], robotics [5], and geoscience [41]. For example, the generative modeling of a protein requires descriptors of its position, orientation, and torsion angles for each amino acid. While the position is Euclidean, the orientation lies in the 3D rotation group $\mathrm{SO}(3)$ and the torsion angle lies in the flat torus where a translation by $2\pi$ will result in the same angle. Existing works for protein design typical require 200 to 1000 sampling steps, limiting the overall throughput for virtual screening. A fast and effective few-step generative model on Riemannian manifolds, therefore, can significantly accelerate the drug discovery and enzyme design processes, further facilitating crucial real-world pharmaceutical applications.
While the current Euclidean consistency model has achieved superior performance on images where the distance between two predictions is measured with the standard Euclidean norm, its counterpart in Riemannian manifolds, the Riemannian consistency model (RCM), poses additional challenges and remains largely unexplored. Specifically, the intrinsically curved manifold requires the consistency parameterizations to lie on the manifold. In this way, simple linear interpolation like Lu and Song [33], Yang et al. [54] is not always feasible. Furthermore, such a manifold constraint will impose an additional constraint on the vector field when the consistency loss is enforced, that is, the vector field at different points should also lie in their corresponding tangent spaces, necessitating additional corrections in the time derivative.
@@ 6 Conclusion
Acknowledgement
This work was funded by the DOE Center for Advanced Bioenergy and Bioproducts Innovation (U.S. Department of Energy, Office of Science, Biological and Environmental Research Program under Award Number DE-SC0018420).
@@ B.2 Spherical Manifold
$\exp_{x}(u+\mathrm{d}u)-\exp_{x}u$ $\approx J_{R}^{-1}(y)J_{R}(u)\mathrm{d}u,y=\exp_{x}u,$ (56)
$\exp_{x+\mathrm{d}x}(u)-\exp_{x}u$ $\approx J_{L}^{-1}(y)J_{L}(x)\mathrm{d}x,y=\exp_{x}u.$ (57)
Combining approximations in Eq.55 and 56, we have
@@
The calculation of $\mathrm{d}(\exp u)_{x}$ is more tricky, as one need to combine Eq.54, 55, and 57. The direct perturbation term can be calculated as:
$J_{R}(y)J_{L}^{-1}(y)J_{L}(x)J_{R}^{-1}(x)\mathop{}\!\mathrm{d}x$ $=J_{R}(y)\left(J_{R}^{-1}(y)R_{y}^{\top}\right)\left(R_{x}J_{R}(x)\right)J_{R}^{-1}(x)\mathop{}\!\mathrm{d}x$ (59)
$=R_{y}^{\top}R_{x}\mathop{}\!\mathrm{d}x=(R_{u}^{\top}R_{x}^{\top})R_{x}\mathrm{d}x=R_{u}^{\top}\mathop{}\!\mathrm{d}x,$ (60)
where $R_{u}$ denotes the rotation by the rotation vector $u$. Therefore, combine the covariant derivative term, we obtain:
$\mathrm{d}(\exp u)_{x}(v)=R_{u}^{-1}v-\frac{1}{2}J_{R}(u)[v,u],$ (61)
The time derivative of the model using the Jacobian-vector product (JVP) should also be modified as:
@@
where the additional left inverse Jacobian comes from the expansion of the reference vector field $\dot{x}$ using the approximation relation in Eq.54.
@@ C.3 Training Technique
To further stabilize the gradient variance in consistency models, we build upon the concept of tangent normalization proposed in Lu and Song [33]. We observed that when the tangent norm is small, standard normalization preserves only directional information, which can impede further optimization. Therefore, drawing inspiration from gradient clipping, we introduce a tangent clipping mechanism, detailed below.
@@
where $M$ represents the maximum tangent norm, which we set to 10 in our experiments. This approach is designed to maintain low gradient variance without inappropriately rescaling the tangent when its norm is small.